Solve 9x^4-21x^2+10 = 0
Give an exact answer, using radicals as needed.
I think to start this problem off I need to factor but I'm not sure how to do that on this equation. Please help.

Answer by scott8148(6628)

factoring ___ (3x^2 - 5)(3x^2 - 2) = 0

3x^2 - 5 = 0 ___ 3x^2 = 5 ___ x^2 = 5/3 ___ x = ± sqrt(5/3)

3x^2 - 2 = 0 ___ 3x^2 = 2 ___ x^2 = 2/3 ___ x = ± sqrt(2/3)